OpenStudy (legomyego180):

Why is the limit of n^(1/n) as n approaches infinity 1?

OpenStudy (xfdgbh536t):

3,10,17,24, ... what will be the 100th number in the pattern?

OpenStudy (xfdgbh536t):

3,10,17,24, ... what will be the 100th number in the pattern?

zepdrix (zepdrix):

Well we're approaching this weird indeterminate form \(\large\rm \infty^0\). It's sort of hard to tell which is happening faster though, right? Is the exponent approaching 0 fast enough to kill the base? We can use that old exponent log trick, leading into L'Hopital's Rule to determine what's going on.

zepdrix (zepdrix):

Recall this little trick,\[\large\rm e^{\ln\left(\color{orangered}{x}\right)}=\color{orangered}{x}\]Exponential base e and log base e are inverse operations of one another, so taking their composition gives us the argument back. We want to apply this trick in reverse,\[\large\rm \color{orangered}{\lim_{n\to\infty}n^{1/n}}=e^{\ln\left(\color{orangered}{\lim_{n\to\infty}n^{1/n}}\right)}\]

zepdrix (zepdrix):

Pass the log into the limit using limit laws,\[\Large\rm e^{\lim_{n\to\infty}\ln\left[n^{1/n}\right]}\]And from there, let's ignore the base for now. Just pay attention to what's going on up in the exponent.\[\large\rm \lim_{n\to\infty}\ln\left[n^{1/n}\right]\]

zepdrix (zepdrix):

Log rule allows us to bring the 1/n out in front of the log, ya? And then some stuff from there.

OpenStudy (legomyego180):

man this calc II stuff is killer

OpenStudy (legomyego180):

they werent kidding when they said this would be a tough class. I see what your saying, thanks for explaining it so thoughtfully zep!

zepdrix (zepdrix):

Do you understand how to finish it up from there? :O

OpenStudy (karim728):